OpenStudy (anonymous):

Dan poured 5 cups of water into a rectangular container measuring 8 inches by 6 inches by 7 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

OpenStudy (anonymous):

A. the volume of the container is 8*2.54*6*2.54*7=2167.74 cm^3=2.17dm^3 (all the measures were converted from inch to cm) in 2.17dm^3 you can pour 2.17l of water or 8*6*7/2.54=132.28 cubic inches B. it was poured 14.44*5=72;2 cubic inches into the container C to see the heigth you must divide this volume by the area of the container which is 8*6=48 heigth=72.2/48=1.5 inch=3.82 cm

OpenStudy (anonymous):

I Know a, and B, But I'm lost on how to figure out c :(

OpenStudy (anonymous):

THANKS :D

OpenStudy (anonymous):

My pleasure